# On the Age of Information in Single-Server Queues with Aged Updates

### v1

@@ 4.4.1 Retrial queueing model example
Consider a M/M/1 retrial queue with single re-attempts [^2] In contrast to the model studied in [18, Section 3.5.1], where multiple retrials occur.. Primary arrivals and retrials follow Poisson processes with rates $\lambda$ and $\theta$, respectively. Any primary arrival that finds a busy server joins the orbit. Similarly, if an update attempts to re-enter the system from the orbit and finds the server busy, it returns to the orbit. When the server is idle, the first arrival —whether it is a primary arrival or a retry from the orbit— enters service immediately. The processed updates are delivered at exponential times of rate $\mu$.
This system can be modelled using a Continuous Time Markov Chain with the state space $\left\{(i,n),i\in\{0,1\},\ n\in\mathbb{Z}^{+}\right\}$, where $i$ represents the server state and $n$ is the number of updates in orbit. Figure 7 shows the transitions and rates for this Markov Chain.
Figure 7: Flow diagram for the retrial queue with central dispatcher. Arrivals at the server occur at rates $\lambda$ –fresh, from the source– and $\theta$ –aged, from the orbit–, and departures from the server occur at exponential times of rate $\mu$.
@@
Notably, the probability of finding the server busy in this system proves to be independent of the retrial rate $\theta$, and the same as in a standard M/M/1$\infty$ model:
@@
Moreover, in the limiting case of instantaneous retrials (i.e., $\theta\to\infty$ and, therefore, $\pi\to 1$) the steady-state probabilities in (15) approach those of the M/M/1$\infty$ queue. In that case, the orbit behaves like an ordinary queue, and the random service discipline in the orbit does not affect the long-term state of the system. It is also worth mentioning that, for the probabilities of Theorem 3 to make sense, it is necessary that $\rho<\pi$, which is stronger than the usual stability condition $\rho<1$. Intuitively, this inequality is equivalent to requiring that the retrial rate is high enough to ensure that the orbit does not grow infinitely.
@@
In turn, $W_{o}$ follows from Little’s Law [19] applied to the orbit,
@@ 4.4.2 AAoI using the aged-updates scheme
Figure 8a) shows an example of sample path for the age in the M/M/1 retrial queue. Arrivals at the server, represented by black triangles, occur at epochs $t_{n}$, while departures are marked with inverted black triangles at epochs $t^{\prime}_{n}$. Primary arrivals go directly to the server, while updates arriving at epochs $t^{o}_{n}$ find the server busy and join the orbit, starting service after possibly several reattempts –denoted by crosses in Figure 8(a)–. After departures, the shaded intervals represent the period during which the server remains idle until the next arrival.
@@
The first term in the right hand side of (18) is the AAoI in a M/M/1/1 queue without pre-emption in service. According to Equation (21) in [6], this expression yields
@@
The age process $A_{n}$ correspond to the waiting time in orbit of the incoming updates, which is zero for arrivals from the source and positive for arrivals from the orbit. The random service discipline also ensures the independence between waiting times and future inter-departure times, resulting in the simplified case of Section 4.1. Using (17), the average initial age is then given by
@@ 5 Conclusions
In this work, we relax the idealised assumption that all updates enter the system with zero age, and develop a more general framework to account for scenarios where updates may already be aged upon arrival. This approach models a situation that arises naturally in multi-hop networks, sensor systems with preprocessing delays, and retransmission protocols.
Our main result is a closed-form expression for the AAoI in single-server queues with aged updates. We demonstrate that the AoI for this system is equal to that of a system with zero-age packets, augmented by a correction term that captures the interaction between the packet’s initial age and the inter-departure times. This formulation enables us to recover simpler results under independence and to deduce analytical bounds when complex dependencies are present. We illustrate the applicability of the proposed framework by deriving well-known and new results across a variety of queueing systems, including retransmissions, tandem queues of different complexity, and retrial systems. Additionally, we conduct simulation tests to assess the validity of the theoretical results in a network characterised by a complex dependence structure, highlighting both the generality and practical relevance of the approach.
Several lines for future research emerge from this work. One natural direction is to extend the focus from single-server systems to more complex and realistic scenarios, such as multi-source systems, where aged updates may compete for server resources. Another extension would be to explore the impact of relaxing the assumption of i.i.d. arrivals, where the correlation between queueing times and initial ages becomes more confusing. It could also be interesting to investigate optimal scheduling policies that explicitly account for the initial age of updates. The analysis could also be extended to consider negative initial ages, which may arise in systems where updates remain valid for a certain period before they begin to age. This approach would require revisiting the graphical method used in the proof of the main result, as the current one assumes non-negative initial ages. Another promising research direction is the application of the aged-updates framework to alternative age metrics, including the Age of Incorrect Information or the Query AoI. Finally, it could be valuable to justify the theoretical findings empirically by measuring the age on real data, assessing the practical accuracy of the expressions and bounds derived in this work.
@@ Appendix A Proof of Lemma 1
Let $X\sim\textsl{exp}\left(\lambda\right)$ and $Y\sim\textsl{hypoexp}\left(\mu_{1},\mu_{2}\right)$ independent random variables, $Z=Y-X$, and define the event $\Lambda=\left\{Y>X\right\}$. If
@@
With a little algebra, this expression can be simplified to that of Lemma $\qquad{}_{\square}$
@@ Appendix B Proof of Proposition 1
This queue is a particular case of the G/M/1$\infty$ queue, which has been extensively studied [20, 21]. There are two main consequences of not having Poisson Process-type inputs. First, the PASTA (Poisson Arrivals See Times Averages) property is no longer available: the probability that an arrival sees the system in state $q$ is not the same as the stationary probability of the system being in state $q$. Second, since the inter-arrival times do not have the memoryless property, the residual inter-arrival times (which are also referred to here as idle times, since the server is idle during this period) and the real inter-arrival times do not have the same distribution. Finding the probability distribution of the residual inter-arrival times is another matter of interest.
Losing the PASTA property means that, to study this system, not only the number of packets is needed, but also the time elapsed since the last arrival. To overcome this problem, the system is observed at very specific instants. Let $L_{k}$ be the number of packets in the system just before the $k$th arrival. Then the stationary number of jobs in the system before arrival is $L=\lim_{k\to\infty}L_{k}$ and, by [20], Equation (6.27),
@@
Arguing as in the M/M/1$\infty$ queue,it can be concluded that the system times are exponentially distributed with rate ($\mu-\mu\sigma$), $T_{n}\sim\textsl{exp}\left(\mu-\mu\sigma\right)$. Therefore, Lemma 2 gives the probability
@@
which, by Equation (22a), is equal to $\sigma$. This result can be better understood by realizing that the event $\left\{X_{n}>T_{n-1}\right\}$ is the same as the event {the nth arrival finds an empty $\left\{\text{the $n$th arrival finds an empty system}\right\}$, and that the probability of the latter is $1-\sigma$ by the geometrical distribution (21).
@@
yields Proposition $\qquad{}_{\square}$
Appendix C Proof of Theorem 3
The steady-state solutions for this system can be obtained following the same procedure as in [18, Section 3.5.1]. Let $p_{i,n}$ be the probability that the system is in state $(i,n)$. From the flow diagram of Figure 7, we can derive the balance equations for this system:

### v2

@@ 4.4.1 Retrial queueing model example
Consider a M/M/1 retrial queue with single re-attempts [^2] In contrast to the model studied in [gross:fqt, Section 3.5.1], where multiple retrials occur.. Primary arrivals and retrials follow Poisson processes with rates $\lambda$ and $\theta$, respectively. Any primary arrival that finds a busy server joins the orbit. Similarly, if an update attempts to re-enter the system from the orbit and finds the server busy, it returns to the orbit. When the server is idle, the first arrival —whether it is a primary arrival or a retry from the orbit— enters service immediately. The processed updates are delivered at exponential times of rate $\mu$.
This system can be modelled using a Continuous Time Markov Chain with state space $\left\{(i,n),i\in\{0,1\},\ n\in\mathbb{Z}^{+}\right\}$, where $i$ represents the server state and $n$ is the number of updates in orbit. Figure 7 shows the transitions and rates for this Markov Chain.
Figure 7: Flow diagram for the retrial queue with single retrials. Arrivals at the server occur at rates $\lambda$ –fresh, from the source– or $\theta$ –aged, from the orbit–, and departures from the server occur at exponential times of rate $\mu$.
@@
[…]
@@
[…]
@@
In turn, $W_{o}$ follows from Little’s Law [little:1961] applied to the orbit,
@@ 4.4.2 AAoI using the aged-updates scheme
Figure 8a) shows an example of a sample path for the age in the M/M/1 retrial queue. Arrivals at the server, represented by black triangles, occur at epochs $t_{n}$, while departures are marked with inverted black triangles at epochs $t^{\prime}_{n}$. Primary arrivals go directly to the server, while updates arriving at epochs $t^{o}_{n}$ find the server busy and join the orbit, starting service after possibly several re-attempts –denoted by crosses in Figure 8(a)–. After departures, the shaded intervals represent the period during which the server remains idle until the next arrival.
@@
The first term in the right-hand side of (18) is the AAoI in a M/M/1/1 queue without pre-emption in service. By [costa:2016, Equation (21)], this expression is
@@
The age process $A_{n}$ corresponds to the waiting time in orbit of the incoming updates, which is zero for arrivals from the source and positive for arrivals from the orbit. The random service discipline in orbit means that the waiting time of a packet does not influence when it is selected for service. Once a successful retry occurs, the service is exponential and therefore memoryless. These two facts ensure that waiting times are independent of future inter-departure times, resulting in the simplified case presented in subsection 4.1. Using (17), the average initial age is then given by
@@ 5 Conclusions
This work introduced a general framework for analysing the Age of Information (AoI) in systems where updates arrive with a non-zero initial age, a frequent scenario in multi-hop networks, edge computing, and retransmission protocols. By relaxing the standard assumption of zero-age updates, we developed a more realistic model that captures the impact of prior delays.
[…]
Several promising research directions emerge from this work. One natural extension involves applying the framework to multi-source systems where aged updates compete for server resources. Another direction would explore systems with non-i.i.d. arrival processes, where the correlation structure between queueing times and initial ages introduces additional complexity. The framework also provides a foundation for designing and optimising optimal control policies that exploit initial age information, such as scheduling disciplines that prioritise fresher packets or admission control policies that strategically discard stale updates. The analysis could be extended to consider negative initial ages for systems where updates remain valid for a period before ageing, though this would likely require revisiting the graphical method used in our proof. Another promising direction involves applying the aged-updates framework to alternative age metrics like the Age of Incorrect Information or Query AoI. Finally, empirical validation through real-world measurements would help assess the practical accuracy of the derived expressions and bounds.
@@ Appendix A Proof of Lemma 1
Let $X\sim\textsl{exp}\left(\lambda\right)$ and $Y\sim\textsl{hypoexp}\left(\mu_{1},\mu_{2}\right)$ independent random variables, Z=YXZ=Y-X, and define the event $\Lambda=\left\{Y>X\right\}$. If
@@
[…]
@@ Appendix B Proof of Proposition 1
Proof.
Fig. 3.  Error-prone, zero-wait model as a system with aged updates. a) Error-prone, zero-wait standard model. Blue triangles are successful deliveries and generation of a fresh packet, and red triangles ($t_{3}$ and $t_{5}$) mark failed deliveries and subsequent retransmission of the same packet. b) Equivalent zero-wait, error-free model with aged updates. Failed transmissions in the standard model can be thought of as inducing an independent process of initial ages $A_{n}$.
Figure 3 provides visual support for the demonstration. First, we note that the number of retransmissions for a given packet follows a geometric distribution with probability of success $(1-\alpha)$,
$p_{k}\stackrel{{\scriptstyle def}}{{=}}\mathbb{P}\left[\text{k retransmissions}\right]=\alpha^{k}(1-\alpha),\quad k=0,1,2,\ldots$
Consequently, the effective service time in the zero-wait, error-prone model depicted in Figure 3a) is a random sum of $\textsl{exp}\left(\mu\right)$ times, with the number of terms geometric of parameter $\alpha$. This leads to exponential effective service times with rate $\mu(1-\alpha)$.
Moreover, the probability that the $n$-th packet is fresh equals the probability that the previous delivery was successful; therefore
$\mathbb{P}\left[n\text{-th packet fresh}\right]=\mathbb{P}\left[(n-1)\text{-st delivery successful}\right]=(1-\alpha).$
From Figure 3a), we observe that the zero-wait error-prone model produces four types of packets, classified by whether they are fresh or retransmitted, and whether their delivery is successful or not. The initial age of each packet in the corresponding error-free model depends on its type in the original model. Below, we describe each type in the original model and deduce the corresponding initial age in the equivalent model.
  1. i)
    A packet is fresh and successfully delivered with probability $(1-\alpha)^{2}$. These packets have zero initial age in the equivalent error-free model, like packets 1 and 8 in Figure 3a). Hence, $A_{1}=A_{8}=0$.
  2. ii)
    Packets 2 and 6 in Figure 3a) are fresh but have not been delivered, an event that occurs with a probability $\alpha(1-\alpha)$. In the error-free model, this is equivalent to a packet whose delivery does not change the age process, and therefore behaves as if it arrives with the same initial age as the monitor currently has. This age coincides with the system time of the immediately preceding packet, which in this model is also the effective service time. Hence, $A_{2}\sim A_{6}\sim\textsl{exp}\left(\mu(1-\alpha)\right)$.
  3. iii)
    Packets 5 and 7 in Figure 3a) are retransmissions that are correctly delivered, occurring with a probability $\alpha(1-\alpha)$. In the error-free model shown in Figure 3b), this is equivalent to a packet arriving with an initial age equal to the time accumulated in the system by the original update. This is a geometric sum of $\textsl{exp}\left(\mu\right)$ times where the probabilities are given by $\mathbb{P}\left[k\text{ retransmissions}\right]/\mathbb{P}\left[\text{retransmissions}\right]=p_{k}/(1-p_{0})$. This yields $\textsl{exp}\left(\mu(1-\alpha)\right)$ times for $A_{5}$ and $A_{7}$.
  4. iv)
    Forwarded packets that are not delivered, like packets 3 and 4 in Figure 3a), occur with a probability $\alpha^{2}$. The delivery of these packets does not change the age function, and thus, in the error-free model, it is equivalent to delivering a packet that enters the system with the same age as the monitor currently has. This age is the sum of the system time of the last successfully delivered packet $\textsl{exp}\left(\mu(1-\alpha)\right)$, as shown in part ii)– and the time accumulated in the system by the original packet –also $\textsl{exp}\left(\mu(1-\alpha)\right)$, as shown in part iii). Hence, $A_{3}\sim A_{4}\sim\Gamma\left(2,\mu(1-\alpha)\right)$.
Therefore,
$A_{n}=\begin{cases}0&\text{with probability }(1-\alpha)^{2}\\ \textsl{exp}\left(\mu(1-\alpha)\right)&\text{with probability }2\alpha(1-\alpha)\\ \Gamma\left(2,\mu(1-\alpha)\right)&\text{with probability }\alpha^{2}\end{cases},$
and
$\mathbb{E}\left[A_{n}\right]=\dfrac{2\alpha(1-\alpha)}{\mu(1-\alpha)}+\dfrac{2\alpha^{2}}{\mu(1-\alpha)}=\dfrac{2\alpha}{\mu(1-\alpha)}.$
Furthermore, the initial age process is independent of any other variable in the model because it consists of service times in the zero-wait, error-prone model, which are assumed to be independent. Thus, applying Theorem 1 to the equivalent error-free model results in $\Delta_{zw}^{A}=\Delta_{zw}^{0}+\mathbb{E}\left[A_{n}\right]$, where $\Delta_{zw}=2/\mu$ is the AAoI in a zero-wait, error free model with $\textsl{exp}\left(\mu\right)$ service times. This yields
$\Delta_{zw}^{A}=\dfrac{2}{\mu}+\dfrac{2\alpha}{\mu(1-\alpha)}=\dfrac{2}{\mu(1-\alpha)}\qquad_{\square}$
Appendix C Proof of Proposition 2
@@
[…]
Losing the PASTA property means that, to study this system, not only the number of packets is needed, but also the time elapsed since the last arrival. To overcome this problem, the system is observed at very specific instants. Let $L_{k}$ be the number of packets in the system just before the $k$th arrival. Then the stationary number of jobs in the system before arrival is $L=\lim_{k\to\infty}L_{k}$ and, by [kleinrock:qst1], Equation (6.27),
@@
[…]
@@
which, by Equation (22a), is equal to $\sigma$. This result can be better understood by realizing that the event $\left\{X_{n}>T_{n-1}\right\}$ is the same as the event $\left\{\text{the $n$th arrival finds an empty system}\right\}$, and that the probability of the latter is $1-\sigma$ by the geometrical distribution (21).
@@
[…]
Appendix D Proof of Theorem 3
The steady-state solutions for this system can be obtained following the same procedure as in [gross:fqt, Section 3.5.1]. Let $p_{i,n}$ be the probability that the system is in state $(i,n)$. From the flow diagram of Figure 7, we can derive the balance equations for this system: